# Cutoff for the Swendsen–Wang dynamics on the complete graph

### v1

@@ 6 Projection chain
For ease of notation, let $P_{t}$ be the event that both assumptions A1 and A2 hold; that is, $P_{t}=P_{t}(\hat{C})$ for $\hat{C}$ sufficiently large. Let $S_{t}$ be the event that $A_{t}=A^{\prime}_{t}$. Then,
$\Pr\big{(}\cup_{i=1}^{\ell}S_{t+i}\big{)}$ $\geq\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cup_{i=1}^{\ell}S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})$
$=\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\cdot\big{(}1-\Pr(\cap_{i=1}^{\ell}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})\big{)}.$ (19)
For any desired constant $\delta_{2}\in(0,1)$, we have that $\Pr(P_{t+1}|P_{t})\geq\delta_{2}$ by Fact 6.2. Then by a union bound, we have
$\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\geq 1-\ell(1-\delta_{2}).$
(We note that $\delta_{2}$ depends on $\hat{C}$.)
To get an upper bound for that $\Pr(\cap_{i=1}^{\ell}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1})$, observe
$\Pr(\cap_{i=1}^{\ell}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell}|\cap_{i=1}^{\ell-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cap_{i=1}^{\ell-1}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1}).$ (20)
Then, by the Markov property and Lemma 3.4 we have that for some constant $\delta>0$
$\Pr(\neg S_{t+\ell}|\cap_{i=1}^{\ell-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell}|\neg S_{t+\ell-1},P_{t+\ell-1})\leq 1-\delta.$
To bound the other term in (20), note that
$\Pr(\cap_{i=1}^{\ell-1}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell-1}\mid\cap_{i=1}^{\ell-2}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cap_{i=1}^{\ell-2}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1}).$ (21)
Moreover,
$\Pr(\neg S_{t+\ell-1}\mid\cap_{i=1}^{\ell-2}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})$ $=\Pr(\neg S_{t+\ell-1}\mid P_{t+\ell-2},P_{t+\ell-1})=\frac{\Pr(\neg S_{t+\ell-1},P_{t+\ell-1}\mid P_{t+\ell-2})}{\Pr(P_{t+\ell-1}\mid P_{t+\ell-2})}$
$\leq\frac{\Pr(\neg S_{t+\ell-1}\mid P_{t+\ell-2})}{\Pr(P_{t+\ell-1}\mid P_{t+\ell-2})}\leq\frac{1-\delta}{\delta_{2}}.$
where the second to last inequality follows from Lemma 3.4 and the fact that $\Pr(P_{t+\ell-1}|P_{t+\ell-2})\geq\delta_{2}$. Generalizing this process, we observe that for integer $j$ with $1\leq j\leq\ell-1$
$\Pr(\cap_{i=1}^{\ell-j}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1})=\Pr(\neg S_{t+\ell-j}|\cap_{i=1}^{\ell-j-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cap_{i=1}^{\ell-j-1}\neg S_{t+i}|\cap_{i=1}^{\ell}P_{t+i-1}),$
and
$\Pr(\neg S_{t+\ell-j}\mid\cap_{i=1}^{\ell-j-1}\neg S_{t+i},\cap_{i=1}^{\ell}P_{t+i-1})$ $=\Pr(\neg S_{t+\ell-j}\mid\cap_{k=t+\ell-j-1}^{t+\ell-1}P_{k})$
$=\frac{\Pr(\neg S_{t+\ell-j},\cap_{k=t+\ell-j}^{t+\ell-1}P_{k}\mid P_{t+\ell-j-1})}{\Pr(\cap_{k=t+\ell-j}^{t+\ell-1}P_{k}\mid P_{t+\ell-j-1})}$
$=\frac{\Pr(\neg S_{t+\ell-1},P_{t+\ell-1}\mid\cap_{k=t+\ell-j-1}^{t+\ell-2}P_{k})\cdot\Pr(\cap_{k=t+\ell-j}^{t+\ell-2}P_{k}\mid P_{t+\ell-j-1})}{\Pr(\cap_{k=t+\ell-j}^{t+\ell-2}P_{k}\mid P_{t+\ell-j-1})\Pr(P_{t+\ell-1}\mid\cap_{k=t+\ell-j-1}^{t+\ell-2}P_{k})}$
$=\frac{\Pr(\neg S_{t+\ell-1}\cap P_{t+\ell-1}\mid P_{t+\ell-2})}{\Pr(P_{t+\ell-1}\mid P_{t+\ell-2})}$
$\leq\frac{1-\delta}{\delta_{2}}.$
Hence, we obtain that inductively from (21)
$\Pr(\cap_{i=1}^{\ell-1}\neg S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})\leq\Big{(}\frac{1-\delta}{\delta_{2}}\Big{)}^{\ell-1},$
and plugging these bounds into (19)
$\Pr(\cup_{i=1}^{\ell}S_{t+i})$ $\geq\Pr(\cap_{i=1}^{\ell}P_{t+i-1})\cdot\Pr(\cup_{i=1}^{\ell}S_{t+i}\mid\cap_{i=1}^{\ell}P_{t+i-1})$
$\geq(1-\ell(1-\delta_{2}))\Big{(}1-(1-\delta)\Big{(}\frac{1-\delta}{\delta_{2}}\Big{)}^{\ell-1}\Big{)}\geq\frac{3}{4}$
for a suitable choice of $\hat{C}$, $\delta_{2}$, and $\ell$. ∎
@@ Appendix A Proof of facts about the drift function $F$
[…]
@@
Let $A\sim\pi$. Then, there exist constants $\gamma_{0},\gamma_{1}>0$ such that $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ and $\sum_{i\geq 2}L_{i}(A)^{2}\leq\gamma_{1}n$ with probability at least 0.90.90.9.
@@
[…]
$\mathrm{E}[\hat{M}]$ $\leq n^{2}\Pr(R\not\in J)+\sum_{r\in J}\mathrm{E}[M\mid R=r]\Pr(R=r)$
@@
Finally, the local limit theorem in Theorem 14 from [LL06] implies that for a suitable constant $\gamma_{0}>0$, we have $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ with probability at least 0.950.950.95; the result then follows by a union bound. ∎
We now use this lemma to proof Lemma 4.4 using the well-known connection between the random-cluster and the ferromagnetic Potts distributions. In particular, if $A\sim\pi$ and we perform a coloring step of the SW dynamics, the resulting configuration has distribution $\mu$ [ES88].
@@
and for a suitable $\ell>0$ we can bound this probability by 0.10.10.1. Since by Lemma B.1, $\Pr(\mathcal{A})\geq 0.9$, we obtain that

### v2

@@ 6 Projection chain
We consider the coupling that in each step aims to couple two instances of the projection chain using the coupling from Lemma 3.4. If it succeeds, the two instances remain coupled in all future steps. For ease of notation, let $\mathcal{P}$ be the set of pairs of configurations of the projection chain that satisfy A1 and A2 for a constant $\hat{C}$ sufficiently large, let $\mathcal{S}_{t}$ be the event that $A_{t}=A^{\prime}_{t}$, and let $Z_{t}=(A_{t},A_{t}^{\prime})$. Let $\hat{\mathcal{P}}\subseteq\mathcal{P}$ be the subset of $\mathcal{P}$ that contains pairs of distinct configurations: i.e., the pairs $(A,A^{\prime})\in\mathcal{P}$ such that $A\neq A^{\prime}$.
It suffices to show that for a suitable constant $\ell$, we have $\Pr\big(\cap_{i=1}^{\ell}\neg\mathcal{S}_{t+i}\big)<1/4$. For this, we obtain first the following recurrence:
$\Pr\big(\cap_{i=1}^{\ell}\neg\mathcal{S}_{t+i}\big)=\Pr\big(\neg\mathcal{S}_{t+\ell}\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)$
$\leq\sum_{Z\in\mathcal{P}}\Pr\big(\neg\mathcal{S}_{t+\ell},Z_{t+\ell-1}=Z\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+\Pr(Z_{t+\ell-1}\not\in\mathcal{P})$
$=\sum_{Z\in\hat{\mathcal{P}}}\Pr\big(\neg\mathcal{S}_{t+\ell},Z_{t+\ell-1}=Z\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+\Pr(Z_{t+\ell-1}\not\in\mathcal{P}),$ (23)
where the first inequality follows from the law of total probability, and the second equality from the fact that under our coupling the probability of $\neg\mathcal{S}_{t+\ell}$ is 0 when $Z\in\mathcal{P}\setminus\hat{\mathcal{P}}$. The right-hand-side of (23) satisfies:
$=\sum_{Z\in\hat{\mathcal{P}}}\Pr\big(\neg\mathcal{S}_{t+\ell}\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i},Z_{t+\ell-1}=Z\big)\Pr(Z_{t+\ell-1}=Z\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i})\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+\Pr(Z_{t+\ell-1}\not\in\mathcal{P})$
$=\sum_{Z\in\hat{\mathcal{P}}}\Pr\big(\neg\mathcal{S}_{t+\ell}\mid Z_{t+\ell-1}=Z\big)\Pr(Z_{t+\ell-1}=Z\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i})\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+\Pr(Z_{t+\ell-1}\not\in\mathcal{P})$
$\leq(1-\delta)\sum_{Z\in\hat{\mathcal{P}}}\Pr(Z_{t+\ell-1}=Z\mid\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i})\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+\Pr(Z_{t+\ell-1}\not\in\mathcal{P})$
$\leq(1-\delta)\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+\Pr(Z_{t+\ell-1}\not\in\mathcal{P}),$
where the second equality follows from the Markov property; in particular, note that $\{Z_{t}\}$ is a Markov chain and $Z\in\hat{\mathcal{P}}$ is such that $\neg\mathcal{S}_{t+\ell-1}$ holds and can be dropped from the conditioning. The first inequality follows from Lemma 3.4.
Observe also that for $j\in\{1,\dots,\ell-1\}$, it follows from Fact 6.2 that there is a constant $\delta_{2}>0$ that we can choose arbitrarily close to 1 by taking $\hat{C}$ large enough so that
$\Pr(Z_{t+j}\in\mathcal{P})$ $\geq\sum_{Z\in\mathcal{P}}\Pr(Z_{t+j}\in\mathcal{P}\mid Z_{t+j-1}=Z)\Pr(Z_{t+j-1}=Z)$
$\geq\sum_{Z\in\mathcal{P}}\delta_{2}\Pr(Z_{t+j-1}=Z)\geq\delta_{2}\Pr(Z_{t+j-1}\in\mathcal{P}),$
and iterating we get $\Pr(Z_{t+j}\in\mathcal{P})\geq\delta_{2}^{j},$ since by assumption $Z_{t}\in\mathcal{P}$. Then,
$\Pr\big(\cap_{i=1}^{\ell}\neg\mathcal{S}_{t+i}\big)$ $\leq(1-\delta)\Pr\big(\cap_{i=1}^{\ell-1}\neg\mathcal{S}_{t+i}\big)+1-\delta_{2}^{\ell-1}.$
Iterating once again,
$\Pr\big(\cap_{i=1}^{\ell}\neg\mathcal{S}_{t+i}\big)$ $\leq(1-\delta)^{\ell}+\sum_{j=1}^{\ell-1}(1-\delta_{2}^{j})(1-\delta)^{\ell-j-1}\leq(1-\delta)^{\ell}+(1-\delta)^{\ell}\sum_{j=1}^{\ell-1}\frac{1-\delta_{2}^{j}}{(1-\delta)^{j+1}}.$
Now first fixing $\ell$ so that $(1-\delta)^{\ell}<1/8$, we can then pick $\delta_{2}$ close enough to $1$ so that $\sum_{j=1}^{\ell-1}\frac{1-\delta_{2}^{j}}{(1-\delta)^{j+1}}<1$, and the result follows. ∎
@@ Appendix A Proof of facts about the drift function $F$
where $c(V,A)$ is the number of connected components in the subgraph $(V,A)$ and $Z_{\textsc{rc}}$ is the corresponding partition function. For $A\subseteq E$, let $L_{i}(A)$ denote the size of the $i$-th largest component of $A$ (breaking ties arbitrarily) and let $\theta=\frac{aq-1}{q-1}$.
@@
Let $A\sim\pi$. Then, there exist constants $\gamma_{0},\gamma_{1}>0$ such that $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ and $\sum_{i\geq 2}L_{i}(A)^{2}\leq\gamma_{1}n$ with probability at least $0.9$.
@@
Let $M=\sum_{i\geq 2}L_{i}(A)^{2}$. Consider the process of coloring the components of $A$ red each with probability $1/q$ independently. Let $R$ correspond to the number of red vertices (excluding the largest component of $A$) and $\hat{M}=\sum_{i\geq 2}L_{i}(A)^{2}\xi_{i}$ where the $\xi_{i}$’s are independent $\mathrm{Bernoulli}(1/q)$ random variables. Observe that $\mathrm{E}[M]/q=\mathrm{E}[\hat{M}$]. We derive a bound for $\mathrm{E}[\hat{M}$] and obtain the probabilistic bound for $M$ via Markov’s inequality. Let $J=[\frac{1-\theta}{q}n-\delta n^{8/9},\frac{1-\theta}{q}n+\delta n^{8/9}]$ for a suitably large constant $\delta>0$. Then,
$\mathrm{E}[\hat{M}$] $\leq n^{2}\Pr(R\not\in J)+\sum_{r\in J}\mathrm{E}[\hat{M}\mid R=r]\Pr(R=r)$
@@
Finally, the local limit theorem in Theorem 14 from [LL06] implies that for a suitable constant $\gamma_{0}>0$, we have $|L_{1}(A)-\theta n|\leq\gamma_{0}\sqrt{n}$ with probability at least $0.95$; the result then follows by a union bound. ∎
We now use this lemma to prove Lemma 4.4 using the well-known connection between the random-cluster and the ferromagnetic Potts distributions. In particular, if $A\sim\pi$ and we perform a coloring step of the SW dynamics, the resulting configuration has distribution $\mu$ [ES88].
@@
and for a suitable $\ell>0$ we can bound this probability by $0.1$. Since by Lemma B.1, $\Pr(\mathcal{A})\geq 0.9$, we obtain that